OpenStudy (anonymous):

sin^-1(sin9pi/7) can someone help me by explaining how I can find the angle that lies in the interval [-2/pi,2/pi] to get the correct answer?

Parth (parthkohli):

Note that \(x \) is a solution for \(\sin^{-1}(\sin (x))\), but it's not the only solution for obvious reasons.

Parth (parthkohli):

\(x + 2\pi\), \(x - 2\pi\) are solutions too.

Parth (parthkohli):

So \(9\pi/7\) is a solution.

OpenStudy (anonymous):

thank you for the explanation. what if it was [-pi/2,pi/2] ? what would that change?

Parth (parthkohli):

Remember the graph of sine. That always helps ya.